# First Light and Reionization Epoch Simulations (FLARES) - XV: The physical properties of super-massive black holes and their impact on galaxies in the early Universe

### v2

@@ 4 Relation to Host Galaxy Physical Properties
We now explore the correlations between SMBH properties and their hosts.
@@ 4.1 $M_{\bullet}-M_{\star}$ scaling relationship
Figure 7: The relationship between stellar mass ($M_{\star}$) and SMBH mass ($M_{\bullet}$) predicted by FLARES for galaxies at $z=5$. In the top-panel only FLARES predictions are shown with individual objects are colour-coded by their Eddington ratio $\lambda$. The diagonal lines denote fixed values of $M_{\bullet}/M_{\star}$. The lower-panel instead shows $M_{\bullet}/M_{\star}$ across $105<z<$ 10, but restricts to objects with $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$, and also includes observations at $z\sim 5$ from Maiolino et al. (2023b), Harikane et al. (2023) and Kocevski et al. (2023).
@@ 4.1.1 Comparison with Observational Constraints
The top-panel of Figure 7 shows a comparison of the FLARES predictions with the local empirical relations of Kormendy & Ho (2013b) and Reines & Volonteri (2015). If restricted to resolved SMBHs the FLARES predictions closely match the Kormendy & Ho (2013b) trend but lie somewhat above the Reines & Volonteri (2015) relation.
With the advent of JWST, measurements have become possible to higher redshift and lower luminosity. Several studies (e.g. Kocevski et al., 2023; Harikane et al., 2023; Maiolino et al., 2023b) have now measured stellar and SMBH masses of galaxies at $z>5$, albeit with small samples. These observations are included in the lower-panel of Figure 7 where we limit the FLARES predictions to $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$ to better align with the completeness of the observations (and also present $M_{\bullet}/M_{\star}$ instead of $M_{\bullet}$ on the $y$-axis). This reveals an overall good correspondence with the majority of observational constraints intersecting with the FLARES predictions. It is worth noting that the observational constraints extend to $M_{\star}<10^{9}\ {\rm M_{\odot}}$, and where they do, the $M_{\bullet}/M_{\star}$ ratios are high (see Maiolino et al., 2023b). Such objects are not predicted by FLARES due to the mass resolution.
@@ 4.2 Relative Contribution of SMBHs to the Bolometric Luminosities of Galaxies
Figure 9: The fraction of galaxies in which $L_{\bullet}>L_{\star}$ (solid thick lines) and $L_{\bullet}>0.1\times L_{\star}$ (dashed thin lines) as a function of the total bolometric luminosity for $z=5\to 10$.
We next explore the relative contribution of SMBH emission to the total bolometric luminosity of galaxies. In Figure 8 we present the ratio of the SMBH to stellar bolometric luminosity as a function of the total bolometric luminosity. The bolometric luminosities of objects with ratios $>1$ are then dominated by emission from a SMBH. In Figure 9 we instead show the fraction of galaxies in which the SMBH provides $>50\%$ or $\gtrsim 10\%$ of the bolometric luminosity.
While there is a large amount of scatter in this relation there is a clear trend of an increase in the $L_{\bullet}/L_{\star}$ ratio with total bolometric luminosity. Objects with $L_{\rm bol}>10^{45.5}\ {\rm erg/s}$ on average have SMBH bolometric luminosities surpassing starlight. Below $L_{\rm bol}\sim 10^{45}\ {\rm erg/s}$ the average fractional contribution appears to drop significantly. However, this may reflect the incompleteness in the black hole population in FLARES due to the resolution and modelling choices. From Figure 9 it is also clear that the relative contribution of SMBHs at fixed total bolometric luminosity drops with decreasing redshift. For galaxies with $L\sim 10^{45.5}\ {\rm erg\ s^{-1}}$ virtually all objects are SMBH dominated at $z=10$, dropping to $\approx 30\%$ at $z=5$.
This means that FLARES, unlike EAGLE, is exploring the regime in which emission from SMBHs is dominating the bolometric emission in galaxies. However, FLARES is only just doing this, and only simulates a handful of SMBHs with bolometric luminosities $>10^{46}\ \rm erg\ s^{-1}\approx 2\times 10^{12}\ {\rm L_{\odot}}$.
@@ 4.3 Total Bolometric Luminosity Function
Building on this analysis, in Figure 10 we show the bolometric function of stars, SMBHs, and the total at $105<z<$ 10. As anticipated from Figure 8 the bolometric luminosity function is dominated by stars at $L_{\rm bol}<10^{45}\ {\rm erg/s}$ but becomes dominated by SMBHs at $L_{\rm bol}>10^{45.5}\ {\rm erg/s}$, providing a clear boost to the total bolometric luminosity function.
@@ 5 Conclusions
In this work, we have explored the physical and photometric properties of super-massive black holes (SMBHs) at high-redshift ($5\leq z\leq 10$) using the First Light And Reionisation Epoch Simulations (FLARES). FLARES is a suite of hydrodynamical zoom simulations probing a wide range of environments $\delta_{14}\approx-0.3\to 0.3$ making it ideally suited to studying rare objects. FLARES employs the well tested EAGLE physics model with the novel re-simulation strategy and post-processing to predict observational quantities. We predict the physical properties of SMBHs and their host galaxies and, through simulations without AGN feedback, we explore the effect of AGN on their host galaxies. Our conclusions are:
  • The resolution and modelling choices of FLARES limit us to the study of SMBHs with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. Compared to the $(100\ {\rm Mpc})^{3}$ EAGLE reference volume, FLARES simulates $\approx 8-20$ times more $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs at $z=5-9$ and $\approx 25$ times more $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs at $z=6$, with samples of $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs extending to $z=9$.
@@
  • The number density of SMBHs predicted by FLARES drops by 10$\approx$ 10$\times$ 10 × from $z=5\to 7$. This trend may continue to higher-redshift but beyond $z=8$ the number of SMBHs simulated by FLARES is small. The density of SMBHs also drops by 10$\approx$ 10$\times$ 10 × for SMBHs with $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$. FLARES predictions are compatible with recent observations of the $z=5$ broad line SMBH mass function; however there is significant uncertainty about the required completeness correction required to convert the broad line mass function to a total SMBH mass function. Where they overlap (in mass and redshift) FLARES is in good agreement with some other models, including Bluetides, Illustris, and TNG100, but lies above models including Astrid, Simba, and TNG300.
@@
  • SMBHs are preferentially found in over-dense environments. The densest regions simulated by FLARES ($\delta_{14}=0.3$) have a density of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs 20$\approx$ 20$\times$ 20 × higher than mean density regions. One implication is that observational studies will be strongly affected by cosmic variance.
@@
  • At fixed mass, SMBHs in FLARES exhibit a range of accretion rates, with almost all having $10^{-6}\leq\lambda\leq 1$. The median $\lambda$ is $\approx 10^{-2}$, though decreases with $M_{\bullet}$.
@@
  • The predicted SMBH bolometric luminosity evolves by $\sim 100$ from $z=10\to 5$. At $z=5-6$ it provides a close match to pre-JWST constraints. While the $z=5$ luminosity function is marginally consistent with JWST observations, at higher redshift there is increasing disagreement, particularly at the bright end. Some of this disagreement can be ameliorated by modelling choices, i.e. the choice of radiative efficiency and averaging timescale.
@@
  • The contribution of SMBHs to the bolometric luminosities of galaxies is found to rapidly increase as a function total bolometric luminosity. However, some of this increase may reflect incompleteness due to the resolution and modelling choices. For galaxies with $L_{\rm bol}>10^{45.5}\,{\rm erg\ s^{-1}}$ we find that accretion on to SMBHs, on average, dominates the total bolometric luminosity. This is reflected in the impact of SMBHs to the bolometric luminosity function (LF) of galaxies with SMBHs dominating at $L_{\rm bol}>10^{45.5}\,{\rm erg\ s^{-1}}$.
@@
The advent of JWST has provided new observational tools to study galaxy formation and evolution in the distant Universe, particularly the role and contribution of SMBHs. While we have only begun scratching the surface of JWST’s contributions, there are already indications of tension between these observations and the FLARES predictions suggesting improvements to the physical model are required.
Looking to the future, JWST will continue building up observations - both larger imaging surveys, allowing us to photometrically identify AGN, and spectroscopic follow-up providing the means to unambiguously determine the contribute of AGN in composite objects. Furthermore, the Euclid spacecraft is now embarking on its mission to map almost the entire extragalactic sky in the optical and near-IR. These observations will allow us to identify bright/rare candidate AGN for subsequent follow-up by JWST.
While FLARES extends the range of masses and luminosities probed by hydrodynamical models with similar mass resolution it is still unable to reach the limit of observational studies. A second phase of FLARES is now underway which will both increase the simulated volume and resolution but also explore a wider set of models and model variations.
@@
In Version 1 of this manuscript we used the number densities presented in the original submission of Kokorev et al. (2024) in Figure 10. It was subsequently discovered that, due to an analysis error, these values were erroneously high by $\approx 0.2$ dex. In this version we use updated values from Kokorev et al. (2024). This change improves the agreement between our predictions and Kokorev et al. (2024), though a discrepancy still exists at $z>5$, particularly in the most luminous bins.
@@ Acknowledgements
We thank the EAGLE  team for their efforts in developing the EAGLE  simulation code. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@
Appendix A Instantaneous vs. time averaged accretion rates
In this work we have utilised the instantaneous accretion rates. However, since, in the FLARES/EAGLE model these rates can vary dramatically between time-steps it is interesting to explore how averaging on various timescales affects some of our results.
We being in Figure 12 by showing the ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 50 Myr for SMBHs at $z=5$. This reveals that the majority of galaxies have average accretion rates larger than their instantaneous rate. Despite this, the total amount of accretion, across all black holes, is actually larger, albeit only slightly ($<10\%$), for instantaneous accretion rates.
Figure 12: The ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 50 Myr for SMBHs at $z=5$. The solid line denoted the median in $M_{\bullet}$ bins. Individual BHs are colour coded by their Eddington ratio ($\lambda$).
This effect has important implications for the accretion rate distribution function, or bolometric luminosity function. Figure 13 shows the bolometric luminosity function calculated assuming the different accretion timescales $\in\{10,20,50,100,200\}\ {\rm Myr}$. While the density of very-luminous ($>10^{46}\ {\rm erg\ s^{-1}}$) SMBHs remains relatively unchanged, the number of lower luminosity, but still bright ($>10^{44.5-45.5}\ {\rm erg\ s^{-1}}$) SMBHs increases significantly. This is due to many BHs with very low instantaneous accretion rates having significantly higher averaged rates as demonstrated by Figure 12.

### v3

@@ 4 Relation to Host Galaxy Physical Properties
We now explore the correlations between SMBH properties and their hosts. In common with the other FLARES analyses, the properties (e.g. the stellar mass - $M_{\star}$) are based on all star particles, associated with the bound sub-halo, within a 30 kpc aperture centred on the potential minimum of the sub-halo.
@@ 4.1 $M_{\bullet}-M_{\star}$ scaling relationship
Figure 9: The relationship between stellar mass ($M_{\star}$) and SMBH mass ($M_{\bullet}$) predicted by FLARES for galaxies at $z=5$. In the top-panel only FLARES predictions are shown with individual objects colour-coded by their Eddington ratio $\lambda$. The diagonal lines denote fixed values of $M_{\bullet}/M_{\star}$. The lower-panel instead shows $M_{\bullet}/M_{\star}$ across $105<z<$ 10, but restricts to objects with $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$, and also includes observations at $z\sim 5$ from Maiolino et al. (2023b), Harikane et al. (2023) and Kocevski et al. (2023).
@@ 4.1.1 Comparison with Observational Constraints
With the advent of JWST, measurements of both SMBH and stellar masses have become possible to redshifts and luminosities probed by FLARES. Several studies (e.g. Kocevski et al., 2023; Harikane et al., 2023; Maiolino et al., 2023b) have now measured stellar and SMBH masses of galaxies at $z>5$, albeit with small samples. These observations are included in the lower-panel of Figure 9 where we limit the FLARES predictions to $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$ to better align with the completeness of the observations (and also present $M_{\bullet}/M_{\star}$ instead of $M_{\bullet}$ on the $y$-axis). This reveals an overall good correspondence with the majority of observational constraints intersecting with the FLARES predictions. However, a small number of the Maiolino et al. (2023b) observations extend to mass ratios beyond those predicted by FLARES at low stellar masses. While the inferred properties of these objects is subject to large uncertainty this may be hinting at a further potential tension with the model. Possibilities here include the resolution of the simulation but also the choice of halo-seeding mass, seed mass, the limitations on growth, and even the feedback. Like the other possible tensions it is clear that further observations are required.
@@ 4.2 Relative Contribution of SMBHs to the Bolometric Luminosities of Galaxies
We next explore the relative contribution of SMBH emission to the total bolometric luminosity of galaxies. In Figure 10 we present the ratio of the SMBH to stellar bolometric luminosity as a function of the total bolometric luminosity. The bolometric luminosities of objects with ratios $>1$ are then dominated by emission from a SMBH.
While there is a large amount of scatter in this relation there is a clear trend of an increase in the $L_{\bullet}/L_{\star}$ ratio with total bolometric luminosity.
Objects with $L_{\rm bol}>10^{46.5}\ {\rm erg/s}$ ($>10^{12.9}\ {\rm L_{\odot}}$) on average have SMBH bolometric luminosities surpassing starlight. At least at $5\leq z\leq 7$ this transition threshold remains stable. This transition luminosity is similar to the $z=3$ transition luminosity in Hopkins et al. (2010) model but falls short of the transition inferred at $z=4-6$, albeit with a large uncertainty. At $z>7$ the transition may be shifting to lower luminosity, however the number of sources, and therefore the significance, is very low. Below $L_{\rm bol}\sim 10^{45.5}\ {\rm erg/s}$ the average fractional contribution drops significantly. This sharp drops like reflects the choice of halo-seeding mass and seed mass in FLARES. It is also worth highlighting that there is a large amount of scatter at fixed total bolometric luminosity. For example, even at $L_{\rm bol}\approx 10^{45}\ {\rm erg/s}$ at $z=5$, where on average (median) SMBHs only contribute $\approx 0.1\%$ of the total bolometric luminosity, around $1\%$ are dominated by their SMBH.
@@ 4.3 Total Bolometric Luminosity Function
Building on this analysis, in Figure 11 we show the bolometric function of stars, SMBHs, and the total at $105<z<$ 10. As anticipated from Figure 10 the bolometric luminosity function is dominated by stars at $L_{\rm bol}<10^{46}\ {\rm erg/s}$ becoming dominated by SMBHs above, at least at $z=5$.
@@ 5 Conclusions
In this work, we have explored the physical and limited photometric properties of super-massive black holes (SMBHs) and their hosts at high-redshift ($5\leq z\leq 10$) using the First Light And Reionisation Epoch Simulations (FLARES). FLARES is a suite of 40 hydrodynamical zoom simulations employing a variant of the EAGLE physics model. The re-simulations encompass a wide range of environments ($\delta_{14}\approx-0.3\to 0.3$), with a bias to extreme over-densities, making it ideally suited to studying rare, massive, and luminous objects. As a consequence of this strategy, despite simulating a volume only slightly larger than the EAGLE reference $(100\ {\rm Mpc})^{3}$ simulation FLARES simulates $\approx 8-20$ times more $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs at $z=5-9$ and $\approx 25$ times more $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs at $z=6$, with samples of $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs extending to $z=9$. By appropriately weighting each simulation we are able to recover, but also crucially extend, key distribution functions and scaling relations. Our conclusions are:
@@
  • The number density of SMBHs predicted by FLARES drops by 10$\approx$ 10$\times$ 10 × from $z=5\to 7$. This trend may continue to higher-redshift but beyond $z=8$ the number of SMBHs simulated by FLARES is small. The density of SMBHs also drops by 10$\approx$ 10$\times$ 10 × for SMBHs with $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$. FLARES predictions are compatible with recent observations of the $z=5$ broad line SMBH mass function; however there is significant uncertainty about the required completeness correction required to convert the broad line mass function to a total SMBH mass function. Where they overlap (in mass and redshift) FLARES is in relatively good agreement with some other models, including Bluetides, Illustris, and TNG100, but lies significantly above models including Astrid, Simba, and TNG300. This reflects the different modelling choices. However, since all the models make multiple changes identifying the decisive factor driving the differences is currently impossible.
@@
  • SMBHs are preferentially found in over-dense environments. The densest regions simulated by FLARES ($\delta_{14}=0.3$) have a density of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs 20$\approx$ 20$\times$ 20 × higher than mean density regions.
@@
  • At fixed mass, SMBHs in FLARES exhibit a range of accretion rates, with almost all having Eddington ratios $10^{-6}\leq\lambda_{\rm Edd}\leq 1$. The median $\lambda_{\rm Edd}$ is $\approx 10^{-1}$, though decreases with $M_{\bullet}$.
@@
  • The predicted normalisation of the SMBH bolometric luminosity function evolves by a factor of $\sim 100$ from $z=10\to 5$. At $z=5-6$ it provides a reasonable match to pre-JWST constraints. While the $z=5$ luminosity function is consistent with JWST observations, at higher redshift there is increasing disagreement, particularly at the bright end with the observational inferred number densities much higher than predicted by FLARES. This tension may have an observational root, for example in the derivation of bolometric luminosities or in the mis-identification of AGN at the highest redshifts. This tension may also reflect modelling choices, for example the choice of radiative efficiency and accretion limit. However, the impact of changing this parameters is complex as enhanced accretion at one epoch may drive a suppression at later times due to increased feedback.
@@
  • The contribution of SMBHs to the bolometric luminosities of galaxies is found to rapidly increase as a function total bolometric luminosity. For galaxies at $75<z<$ 7 with $L_{\rm bol}>10^{46.5}\,{\rm erg\ s^{-1}}$ we find that accretion on to SMBHs, on average, dominates the total bolometric luminosity.
@@
5.1 Future directions
Looking to the future, JWST will continue building up observations - both larger imaging surveys, allowing us to photometrically identify AGN, and spectroscopic follow-up providing the means to unambiguously determine the contribute of AGN in composite objects. Furthermore, the Vera Rubin Observatory and Euclid spacecraft are now embarking on their missions to map much of the extragalactic sky across the optical and near-IR. These observations will allow us to identify bright/rare candidate AGN for subsequent follow-up, and crucially confirmation, by JWST. Synergies with observatories at other wavelengths will increasingly improve constraints on bolometric luminosities, minimising the need to rely on empirical bolometric corrections obtained at low-redshift.
On the modelling side it is evident that FLARES, through its unique strategy, has provided new insights into early SMBH formation and evolution. However, FLARES is only just probing the regime in which AGN dominate the emission of galaxies. To fully exploit wide-area observational surveys it is essential that even larger simulations are conducted, all while maintaining sufficiently high-resolution. Moreover, except for a handful of tests, FLARES adopted a single physics model and parameter set. Ideally, we would systematically vary the key parameters governing SMBH seeding, growth, dynamics, and feedback to gain clearer insights. We are now preparing for a new phase of the FLARES project in which we will not only employ an updated physics model, but simulate even larger effective volumes and with higher resolution. In the longer term we will also carry out ensemble suites, varying all of the key parameters of the model, and even exploring entirely different modelling approaches.
@@ 5.1 Future directions
In Version 1 of this manuscript we used the number densities presented in the original submission of Kokorev et al. (2024) in Figure 11; it was subsequently discovered that, due to an analysis error, these values were erroneously high by $\approx 0.2$ dex. In this version we use updated values from Kokorev et al. (2024) which improves the agreement between our predictions and Kokorev et al. (2024), though a discrepancy still exists at $z>5$, particularly in the most luminous bins. In the process of addressing comments from the two reviewers we also made several other significant changes. Instead of using the instantaneous accretion rates we decided to use the accretion rates averaged over 10 Myr, shifting the discussion of this from the Appendix to the main body of the text. This has a subtle impact on all Figures from Figure 6 on-wards. We also uncovered a bug in the calculation of stellar bolometric luminosities which led them to be under-reported. This has a significant impact on what are now Figures 10 and 11.
@@ Acknowledgements
We would like to thank the two referees for their comprehensive reviews which we believe have resulted in an significantly improved manuscript. We thank the EAGLE  team for their efforts in developing the EAGLE  simulation code. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.